OpenStudy (anonymous):

Solve. 0.4x + 8 <= 1.2x - 5

OpenStudy (katrinakaif):

Bring all the x's to one side by substrating 0.4x from 1.2x 0.4x + 8 <= 1.2x - 5 + 8 <= 1.2x - 0.4 - 5 Combine like Terms (1.2x - 0.4x) 8 <= 0.8x - 5 Bring 5 to the other side by adding it: 8 + 5 <= 0.8x Add 8 + 5 = 13 <= 0.8x Divide by 0.8 13/0.8 < = x 16.25 < = x

OpenStudy (anonymous):

0.4x +8 = 1.2x - 5 transpose 1.2x to left hand side 0.4x-1.2x + 8 = -5 transpose 8 to right hand side 0.4x-1.2x = -5-8 -0.8x = -13 x = -13/(-0.8) x = 13/8 * 10 x = 16.25 Hope u got it

OpenStudy (katrinakaif):

Kush the inequality has to be x is greater than or equal to 16.25.. Equals alone would be incorrect but good try Kush =)

OpenStudy (anonymous):

The answer I got isn't correct so I am doing something wrong.

OpenStudy (anonymous):

Thanks Kat!

OpenStudy (katrinakaif):

Anytime Bunny! =)

OpenStudy (anonymous):

sorry but dumbunny u must post the questions with the help of equation forum use that

OpenStudy (anonymous):

f(x)=-4x-3 find the slope and y -intercept